### v2

# Language-Agnostic Modeling of Source Reliability on Wikipedia

Abstract.
Over the last few years, content verification through reliable sources has become a fundamental need to combat disinformation. Here, we present a language-agnostic model designed to assess the reliability of sources across multiple language editions of Wikipedia. Utilizing editorial activity data, the model evaluates source reliability within different articles of varying controversiality such as Climate Change, COVID-19, History, Media, and Biology topics. Crafting features that express domain usage across articles, the model effectively predicts source reliability, achieving an F1 Macro score of approximately 0.80 for English and other high-resource languages. For mid-resource languages, we achieve 0.65 while the performance of low-resource languages varies; in all cases, the time the domain remains present in the articles (which we dub as permanence) is one of the most predictive features. We highlight the challenge of maintaining consistent model performance across languages of varying resource levels and demonstrate that adapting models from higher-resource languages can improve performance. This work contributes not only to Wikipedia’s efforts in ensuring content verifiability but in ensuring reliability across diverse user-generated content in various language communities.
Wikipedia, knowledge equity, knowledge integrity, information reliability
copyright: nonejournalyear: 2024doi: XXXXXXX.XXXXXXXccs: Information systems Wikisccs: Computing methodologies Machine learning algorithms
1. Introduction
The proliferation of fake news in recent years is deteriorating the reliability and trustworthiness of the online information ecosystem (Del Vicario et al., 2016; Vosoughi et al., 2018). To mitigate their spread and impact in society, researchers on web and data mining are devoting a great deal of effort in generating resources to detect untrue content on social media platforms (Su et al., 2023; Xu et al., 2022; Ashkinaze et al., 2024; Xiao et al., 2024; Guo et al., 2023b; Xiao et al., 2023; Wu et al., 2023).
In popular web platforms like Wikipedia [^1] Wikipedia is among the top visited websites worldwide in 2024: https://www.semrush.com/trending-websites/global/all., disinformation is also a main threat to knowledge integrity (Aragón and Sáez-Trumper, 2021). Wikipedia’s conception of knowledge as a service requires content to be verifiable in order to preserve its reliability and integrity [^2] https://research.wikimedia.org/knowledge-integrity.html. Thus, the platform’s policy on verifiability implies that readers should be able to check that all the information in Wikipedia articles comes from reliable sources [^3] https://en.wikipedia.org/wiki/Wikipedia:Verifiability.. While a classical patrolling technique in this platform is to detect and remove statements that violate basic core content policies (Kumar et al., 2016), some Wikipedia editors have proven more effective in tackling disinformation by first identifying unreliable sources (Cohen, 2021). Recent research revealed the positive impact of the community-curated list of perennial sources in English Wikipedia (Baigutanova et al., 2023a). However, that list is limited or even non-existent in other language editions (Baigutanova et al., 2023b).
To address this challenge, we present our research on source reliability, based on language-agnostic approaches and edit activity data from multiple Wikipedia language editions. We aim to support Wikipedia editors, including the ones in medium and small projects that often miss advanced tools, in identifying sources with patterns associated with low reliability, and in monitoring their prevalence across language editions. To meet this goal, we address the following research questions:
  • RQ1. What are the most predictive language-agnostic features when modeling source reliability in Wikipedia, considering the Climate Change topic in English as a case study?
  • RQ2. To what extent can model performance be related to the topic and size of a language edition of Wikipedia?
  • RQ3. Is it possible to adapt these models across topics and/or languages?
Our methods to answer these questions are inspired by existing language-agnostic approaches to measure the controversiality of linked elements in a given Wikipedia article (Borra et al., 2015). Results with datasets from different topics and language editions suggest a promising scenario. Given the several limitations outlined in the following sections, we pay attention to model performance not only for the largest editions but also in mid- and low-resource settings, where adaptation of the models may provide the most benefit. To our knowledge, this is the first work modeling source reliability on Wikipedia using only language-agnostic features.
2. Related Work
Several research works have assessed the verifiability of Wikipedia content and the reliability of the sources by analysing, e.g., the quality of references, the appropriateness of links between articles, and the consensus between editors regarding this content.
The Contropedia project (Borra et al., 2015), for example, aimed at identifying and visualizing controversial links between Wikipedia articles, i.e., links that point to other articles in the platform and whose collaborative edit history shows signs of dispute between editors before reaching consensus. Furthermore, several tools provide interactive visualizations that help understand the path to consensus for a Wikipedia article  (Borra et al., 2015; Guo et al., 2023a). Cultural differences in the controversiality of specific topics have been shown to exist (e.g., (Fu et al., 2023)).
The Citation Detective tool (Chou et al., 2020) leveraged machine learning models (Redi et al., 2019) to assess the evolution of reference quality between 2010 and 2020, in terms of the ratio of sentences missing a required citation, and the ratio of citations that reference non-authoritative sources. The authors found out that, thanks to the efforts of the community, the ratio of sentences missing a required citation was significantly reduced, while the ratio of references to non-authoritative sources has remained below $1\%$. More recently,  (Petroni et al., 2023) has proposed a neural-network-based system to improve the verifiability of content.
The observed improvements in reference quality coincide with the introduction of the perennial source list (Wikipedia contributors, 2024), a curated, non-exhaustive list of sources whose reliability and use on Wikipedia are frequently discussed. Though the perennial source list has been used by several works to assess the authoritativeness of sources, only a few language versions of Wikipedia contain lists of perennial sources and, as the reliability of a source is not static, these lists might evolve through time. Moreover, (Baigutanova et al., 2023b) showed that the reliability of references might be different across cultures, such that unreliable references in one language might persist across editions in another language. This makes it difficult to train general models for source controversiality and, given that some language editions of Wikipedia are still small, achieving reference quality across all languages calls for language-agnostic approaches as the one presented in this work. Beyond the perennial source list, other domain quality rankings for news sources exist, such as the Iffy index of unreliable news sources  (Barret, 2021) based on the Media Bias/Fact-Check (Zandt, 2024). However, we show that the latter corresponds only loosely with the perennial source list, pointing to the necessity of Wikipedia-specific models.
[…]
The multilingual nature of Wikipedia calls for specific modeling techniques. Models such as multilingual BERT have been used to detect vandalism (Trokhymovych and Saez-Trumper, 2021) or political bias (Swati et al., 2023), while other works have built rankings of relative quality in multiple language versions of the platform (Lewoniewski et al., 2020, 2019). For some other tasks, and due to the limitations found in languages with smaller Wikipedia datasets, language-agnostic approaches have been used, e.g., to assess article quality (Das et al., 2024), or to perform automatic entity-linking (Gerlach et al., 2021) and topic classification (Johnson et al., 2021). Language-agnostic models using features extracted from the network structure or user behavior have also been used in the more general context of detecting misinformation or low-quality information in social media. In this line, Shu et al. (2019) constructed embeddings of users and news and used diffusion models to trace the paths through which fake news propagates, while Zhao et al. (2021) extracted features from the user interaction network to help identify misinformation in online health communities. We show that, indeed, user-behavior related language-agnostic signals are useful for modeling source reliability, and that they can be adapted across languages to improve performance in low-resource settings.
3. Data and Methods
3.1. Perennial Sources
We use the English Wikipedia perennial source list (Wikipedia contributors, 2024) as ground truth for modeling source reliability. It was collected on March 30, 2023. Although the list exists for 12 other languages, we only use the English version as it is by far the most extensive (see (Baigutanova et al., 2023b) for details on coverage). In English, it consists of a collection of web domains in five categories: blacklisted (569), deprecated (125), generally unreliable (178), no consensus (137), and generally reliable (184). To build a binary classifier, we use the generally reliable domains as positive labels, merge the domains blacklisted, deprecated, and generally unreliable into a single unreliable category (i.e., the negative label) and ignore the no consensus domains.
Table 1. Statistics of the topic-language datasets used in the study grouped by high, mid, and low resource languages. All numbers are averages (apart from the number of languages). Perennial Domains refers to the number of domains for the perennial source list in English found in the different datasets.
Topic # Langs # Articles Revs Revs/Article URLs URLs/Article Domains Perennial Domains
High Climate Change 7 $1378$ 372 133372133372\,133372 133 $269$ 58 6715867158\,67158 671 $49$ 15 3631536315\,36315 363 $203$
COVID-19 7 $898$ 233 905233905233\,905233 905 $260$ 60 1686016860\,16860 168 $89$ $8289$ $206$
Biology Sample 7 13 6071360713\,60713 607 586 171586171586\,171586 171 $43$ 59 7635976359\,76359 763 $6$ 11 7131171311\,71311 713 $148$
History Sample 7 $5710$ 624 428624428624\,428624 428 $109$ 52 7025270252\,70252 702 $12$ 12 1211212112\,12112 121 $166$
Media Sample 7 $5577$ 819 541819541819\,541819 541 $146$ 108 583108583108\,583108 583 $22$ 21 2902129021\,29021 290 $259$
Mid Climate Change 36 $395$ 28 3642836428\,36428 364 $71$ $6224$ $19$ $2599$ $94$
COVID-19 37 $272$ 18 5231852318\,52318 523 $68$ $8984$ $39$ $1788$ $102$
Biology 37 $4153$ 83 5998359983\,59983 599 $20$ $8691$ $3$ $2022$ $61$
History 37 $1680$ 83 2588325883\,25883 258 $49$ $6274$ $7$ $2054$ $67$
Media 37 $1018$ 46 9574695746\,95746 957 $46$ $7932$ $10$ $2416$ $123$
Low Climate Change 56 $115$ $3430$ $29$ $1079$ $12$ $570$ $35$
COVID-19 75 $32$ $916$ $28$ $568$ $22$ $242$ $32$
Biology 41 $593$ 21 0962109621\,09621 096 $35$ $1159$ $4$ $460$ $16$
History 55 $511$ 13 5011350113\,50113 501 $26$ $952$ $6$ $406$ $22$
Media 70 $159$ $2262$ $14$ $528$ $7$ $264$ $30$
3.2. Language Groups
As we aim to test the proposed model on a variety of languages, we collect data on all $326$ available language editions of Wikipedia. We then discard all languages which have fewer then two reliable and unreliable sources according to the English perennial source list. Among the remaining ones, we distinguish between low, middle, and high-resource languages by examining the number of “active users”, defined as “registered users who have made at least one edit in the last $30$ days.” [^4] https://meta.wikimedia.org/wiki/List_of_Wikipedias (accessed on 01-03-2024) We consider the top 5% of the languages having the most active users as high-resource, the subsequent 25% as mid-resource, and the remaining 70% as low-resource languages.
3.3. Article Collection
3.3.1. Topic Selection
In this study, we have focused on articles from the following five topic datasets: Climate Change, COVID-19, Biology, History, and Media. Climate Change and COVID-19 were selected as they are both controversial but for a different time duration. COVID-19 appeared at the end of 2019, while Climate Change has been a controversial topic for decades (Korte et al., 2023). Both have a Wikiproject community of editors  [^5] https://en.wikipedia.org/wiki/Wikipedia:WikiProject who label articles as relevant to the topic with special guidelines. For comparison, we also selected three topics that are not as controversial, but which varied in terms of reference quality based on the “reference risk” metric (which is also based on the usage of perennial sources) from previous work (Baigutanova et al., 2023a). We choose one example from each risk category: Biology (low risk), History (medium risk), and Media (high risk).
3.3.2. Data Download
For the first two topics, we use the articles listed by the corresponding topic-based WikiProjects. For the latter three, we use the topics identified using the Wikipedia Objective Revision Evaluation Service (ORES) machine learning system (Halfaker and Geiger, 2020). To keep the sizes of datasets comparable and manageable, we sample the articles identified by ORES to be in Biology and History topics by 10% (each), and in the Media topic by 2%. These two approaches allow us to find articles in the English Wikipedia for each of the five topics, resulting in a total of 84 0278402784\,02784 027 articles. To find the versions of these articles in other languages, we use the MediaWiki API, which results in another 335 809335809335\,809335 809 of articles in all of the 326 languages. We collect the articles and their complete edit history using the MediaWiki API in between December 2023 and February 2024.
3.3.3. Extraction of Source Domains
Citations may appear in the body of the article text and as references at the bottom of the article, each of which has standard templates. Each revision that involves a citation contains the author of the revision, position within the document of the citation, text of the citation, and meta information, such as whether it is a DOI, URL, ISBN, ISSN, etc. or a raw URL. We standardize these revisions to create a list of “source edits” containing the metadata about each insertion, deletion, or edit that involves a source. We resolve redirects, apply rules to clean and standardize the URLs (including DOI references redirecting to a resource) and merge revisions when the same user edits an article in succession (regardless of the time). Finally, we extract the domains from the URLs, obtaining 376 566376566376\,566376 566 unique source domains.
3.3.4. Dataset Statistics
The summary statistics about the topic-language datasets (or simply “datasets” throughout this paper), for the five topics grouped into low-, mid- and high-resource languages can be seen in Table 1. Recall that we only use datasets that have passed the threshold of having at least 2 reliable and 2 unreliable perennial sources.
The smallest topic in terms of the number of articles is COVID-19, while the highest numbers of revisions per article are found around Climate Change and COVID-19 in high-resource languages. On the contrary, biology has the largest number of articles, but these are revised on average the fewest number of times in high- and mid-resource languages. High-resource languages have much better coverage of the English perennial sources, with an average of 150 to 260 sources per topic. These values decrease to an average of 15 to 35 sources for low-resource languages, making the training and evaluation more challenging.
3.4. Feature Definition
Unlike previous studies measuring Wikipedia edit quality that did not consider certain types of edits and reverts (Trokhymovych et al., 2023), we propose to focus on the entire edit history concerning each domain. Using edit metadata, we define 52 features capturing the popularity of the domain, the “permanence” of the domain across edits, and the number and type of users involved in edits adding or removing that domain. As the features are computed for each domain within each dataset separately (recall, by “dataset” we mean a collection of articles on a certain topic, in a certain language) of varying sizes and editorial frequencies, we explore different ways of normalizing these statistics w.r.t. the first appearance of the domain, age of the dataset, and in terms of time duration or number of revisions.
3.4.1. Popularity features
  • $N_{\mathrm{articles}}$, $\overline{N_{articles}}$: Number of articles a domain has appeared in (sensitive to the size of the dataset), and its normalized version (by the number of articles in the dataset).
  • $CurrN_{articles}$, $\overline{CurrN_{articles}}$: Number of articles a domain is used in at collection time, and normalized by the total number of articles in the dataset.
3.4.2. Permanence features
  • $\Sigma Perm_{d}$, $\Sigma Perm_{r}$: Permanence, how long a domain has been used in an article (not necessarily consecutively), summed across all articles and measured in days (subscript $d$) or by this number of revisions (subscript $r$).
  • $\Sigma CurrPerm_{d}$, $\Sigma CurrPerm_{r}$: Same as above, but considering only the articles where the domain is currently used.
  • $\overline{\Sigma Perm_{d}}$, $\overline{\Sigma Perm_{r}}$: Sum of all permanences for a domain in all articles, normalized by the sum of the ages of all articles (measured by the number of days or the number of revisions that an article existed).
  • $\langle Perm_{d}\rangle$, $\langle Perm_{r}\rangle$: The average permanence of a domain (over all articles where it was used).
  • $\langle SelfPerm_{d}\rangle$, $\langle SelfPerm_{r}\rangle$: Self-permanence: permanence divided by the number of days (or revisions) since the first time the domains were added to the article, then averaged article-wise.
  • $\Sigma age_{d}$, $\Sigma age_{r}$: The sum of ages of the domain over all the articles in a dataset, where the age of a domain in an article is the number of days or revisions since a URL from that domain has been first added to the article.
  • $\langle age_{d}\rangle$, $\langle age_{r}\rangle$: The average age of a domain over all the articles it appears at least once.
3.4.3. User-based features
  • $U_{add}$, $U_{rem}$: Number of users that have added or removed a domain.
  • $R_{add}$, $R_{rem}$: Same as above, but only counting registered users.
  • $\overline{U_{add}}$, $\overline{U_{rem}}$, $\overline{R_{add}}$, $\overline{R_{rem}}$: The above four features, normalized by the total number of unique users in the dataset.
  • $\langle U_{add}\rangle$, $\langle U_{rem}\rangle$, $\langle R_{add}\rangle$, $\langle R_{rem}\rangle$: The above four features, measured as an average per article.
  • $Ratio(R_{add}/U_{add})$, $Ratio(R_{rem}/U_{rem})$ The proportion of registered vs. all users that ever added or removed a domain.
  • $Proba(R_{add})$, $Proba(R_{rem})$ Probability that, when a domain is added or removed, this is done by the registered users. Contrary to the ratio here we take into account the number of revisions per user.
  • For all user-based features listed above, we compute a version wherein instead of counting any instance of adding a domain, we only count instances of adding it for the first time on an article (“starting”), and similarly counting only the instances that the domain was removed for the last time (“ending”). For example, $Proba(R_{add})$, and $Proba(R_{rem})$ would have a corresponding version $Proba(R_{start})$ and $Proba(R_{end})$ (similarly for the related features).
When we aggregate the various statistics over articles for a domain, we have considered the average and the median and found the best performance with the average, which we use in the models in the following sections. [^6] We have also experimented with versions of the above features that use the aggregation over all URLs in each dataset, instead of domains. However, the resulting features were performing very similarly to those aggregated over domains, and we exclude these from the analysis..
3.5. Model Training and Evaluation
Using the above features, we train an XGBoost [^7] We use Python package XGBoost V1.6.2 https://pypi.org/project/xgboost/1.6.2/. model to classify the domains as reliable or unreliable. To account for class imbalance, we use weights (e.g., the weight for the positive class is the fraction of negative/positive). To limit overfitting, we set the maximum depth of the classifier to 1 (we have tested the classifier with depths up to 5, but found no increase in leave-one-out (LOO) validation performance, and the learning rate to $\eta=0.1$ (default is 0.3).
When evaluating in the “native” condition, wherein the training and target data are from the same dataset, we employ LOO cross-validation. In the “cross-language adaptation” condition, wherein the model is trained on one language and tested on another (keeping the topic the same), we simply compute the evaluation metrics on the test language. Similarly, for cross-topic adaptation, we keep the language consistent and test on a target topic. In the “mixed” condition, wherein the model is trained on a set of languages and tested on one of these languages, we again employ LOO cross-validation where a domain of test language is left out at a time.
To compute a confidence interval for a metric, we employ bootstrapping wherein we resample with replacement the output of the classifier and recompute the metrics $n=100$ times. The main metric we compute is the F1 macro score, which is the unweighted average of the F1 scores of the negative and positive classes.
4. Results
Table 2. Leave-one-out validation F1 macro, and precision and recall metrics for both classes for the model trained and tested in the English Wikipedia on each topic, and for all topics combined (standard deviations specified with $\pm$). Class balance in terms of the number of reliable (rel) and unreliable (unr) domains.
Topic F1 Macro Precision (rel) Recall (rel) Precision (unr) Recall (unr) # rel # unr
Climate Change $0.81\pm 0.02$ $0.80\pm 0.03$ $0.80\pm 0.04$ $0.83\pm 0.04$ $0.83\pm 0.03$ $129$ 152
COVID-19 $0.88\pm 0.02$ $0.87\pm 0.03$ $0.91\pm 0.03$ $0.89\pm 0.03$ $0.85\pm 0.03$ $136$ 125
Biology $0.80\pm 0.03$ $0.81\pm 0.03$ $0.81\pm 0.03$ $0.80\pm 0.03$ $0.80\pm 0.03$ $124$ 120
History $0.75\pm 0.03$ $0.70\pm 0.04$ $0.83\pm 0.04$ $0.82\pm 0.03$ $0.69\pm 0.05$ $120$ 139
Media $0.83\pm 0.02$ $0.81\pm 0.03$ $0.82\pm 0.03$ $0.85\pm 0.03$ $0.84\pm 0.03$ $146$ 176
All topics $0.83\pm 0.02$ $0.76\pm 0.04$ $0.85\pm 0.03$ $0.89\pm 0.02$ $0.82\pm 0.03$ $159$ 234
Figure 1. Beeswarm analysis of an XGBoost model trained on Climate Change English dataset (class encoding is 1: reliable and 0: unreliable) and two example classification explanations (score shown is the log-odds ratio).
[…]
[¡short description¿]¡long description¿
4.1. RQ1: Case study: Climate Change in English
We begin by studying the behavior of our model for the Wikipedia articles in English around the Climate Change topic. The model performance is shown in the first row of Table 2, which lists the F1 Macro, precision and recall for each class. We find that, out of the unreliable domains, the classifier correctly identifies as unreliable 83% of them (recall), and out of the domains the classifier guessed as unreliable, 83% are truly unreliable (precision). Both precision and recall for the reliable class are slightly worse at 80%. Note that the classes are fairly balanced in this setting, with 46% of the dataset having the reliable label.
Figure LABEL:fig:RQ2_beeswarm shows the “bee swarm” plot of SHAP (SHapley Additive exPlanations) values that illustrate the importance of each feature in the classifier (Lundberg and Lee, 2017). The top 6 most predictive features (and an aggregation of the rest) are shown. The SHAP values for the top feature $Proba(R_{end})$ – the probability that, when a domain is removed, it is done by a registered editor – reveal that high values of this feature are associated with a negative impact in the domain reliability prediction. In other words, when a registered editor removes a domain, it is more likely to be unreliable. The opposite can be seen for the $\langle SelfPerm_{d}\rangle$ –average proportion of days the domain remains cited since it was first added– and $\langle Perm_{r}\rangle$ –the average number of revisions that the domain remains on the article–: higher values of these features are associated with more reliable sources. $\Sigma CurrPerm_{r}$ –the aggregated number of revisions through which the domain has existed across all articles– also correlates positively with the reliable category, so does as well as the sum of the domain ages across articles, $\Sigma age_{r}$. Finally, “younger” domains – those that have been added to an article more recently, $\langle age_{d}\rangle$ – are more likely to be reliable. This may be a sign of the increased focus of the platform on the quality of sources (Baigutanova et al., 2023a).
In Figures LABEL:fig:RQ1_twitter and LABEL:fig:RQ1_Researchgate we illustrate the performance of the model for two domains of special interest: twitter.com, correctly identified as unreliable, and researchgate.net, incorrectly identified as reliable. In the case of twitter.com, despite being added/present frequently on articles (high $\Sigma CurrPerm_{r}$), it is soon removed by registered editors ($Proba(R_{end})$ and also SelfPermanence, $\langle SelfPerm_{d}\rangle$ is low). In the case of researchgate.net, most of the permanence features point to it being used as a reliable source. This points out that, although this domain is considered unreliable in the perennial source list because the platform does not perform fact-checking or peer reviewing and is considered a self-published source, the Wikipedia community often uses this domain as if it were reliable. In fact, this platform could be used as a way to provide an open-access link to peer-reviewed references when the official version is behind a paywall. Therefore, this specific example shows that our methodology could be used to identify cases where the use of domains is not in line with the established guidelines.
4.2. RQ2: Topic & Language Generalizability
We continue by evaluating the performance of our language-agnostic modeling approach in other topical and linguistic settings. Table 2 shows the F1 macro and class-specific precision and recall for these topics, in the English Wikipedia. We find that, overall, the performance is consistent, with the lowest performance observed for the History dataset, with a recall of 0.69 for the unreliable class. The best performance is achieved on the COVID-19 dataset, possibly due to the ample data on the edits and the more recent nature of the corresponding articles. We find that the amount of edited data is important for the classifier’s performance, as we show later in this section. Finally, we combine all English-language datasets into one, and perform training and LOO validation; the performance of this model is shown in Table 2 under “all topics”. Such a combined model performs on par with other topical models, with especially high precision for the unreliable class (0.89), suggesting that combining knowledge about different topics does not overall degrade the performance of the classifier.
Next, we explore the applicability of the proposed model to other languages. In these experiments, we first keep the topic constant – Climate Change – while training and evaluating models in different languages. Figure 2 shows the LOO F1 macro scores for select languages form the high-, mid-, and low-resource groups, in comparison to the performance of a random baseline model. Specifically, we show the average performance $\pm$ one standard deviation as shaded areas. To estimate the performance of a random model, we assign a random binary classification for the domains and perform bootstrapping ($n=100$) similar to the evaluation of all other models, as discussed in Section 3.5. The performance is highest for high-resource languages and degrades for mid- and low-resource ones. The standard deviations of the scores also increase with lower resource languages (both for our model and the baselines), pointing to an increased lack of information for computing the score.
Additionally, we show aggregate statistics of the model performance on the languages in each category in Figure 3. In particular, we are interested in whether our model outperforms the random classifier in terms of F1 macro score (top panel of the figure): we perform this comparison using Mann-Whitney scores and adjust the significance level of 0.05 using the Bonferroni correction for multiple comparisons. For instance, in the Climate Change topic in 100% of the high-resource languages, the models outperform the random baseline, but instead in only 44% of the low-resource languages. The performance on low-resource languages tends to vary widely for different topics. The brown bars show the models for which articles for all topics within a language are used for training and testing. These models perform markedly better than topic-specific ones, echoing our findings for the English language.
Figure 2. Performances (mean $\pm$ stdv) of models trained and validated in languages other than English on the Climate Change datasets. The three batches of languages, separated by a vertical black line, show results on a sample of high (left), mid (centre), and low (right) resource languages.
[…]
[¡short description¿]¡long description¿
Figure 3. Native model performance per topic for different language types. Top panel: % of models that perform better than a random classifier. Bottom panel: Violin plots of corresponding distributions of F1 macro performances.
[…]
[¡short description¿]¡long description¿
Figure 4. Average model F1 macro (line) and standard dev. (shaded area) versus the size of the training dataset in terms of the number of revisions. In English (red), all topics are combined and the data is sampled at regular intervals. The same experiment is repeated, considering only revisions before the implementation of the perennial sources on 2018-07-01 (grey). The performance of models in all other languages (also all-topic) is shown in blue, without sub-sampling.
[…]
[¡short description¿]¡long description¿
Effect of dataset size and introduction of perennial sources on model performance
The worse performance of our model in the lower-resource languages may be mainly (among others) due to two reasons: the datasets available for training the model are too small, or the behaviour of the users of these languages towards domains is intrinsically difficult to model (their treatment of reliable and unreliable domains is difficult to distinguish). We design two experiments to address these two reasons, starting with the model sensitivity to the amount of training data.
The red line in Figure 4 shows the performance of the model trained on different dataset sizes. In particular, we consider the English (all-topic) dataset and sample the data at log-regular intervals (starting at $10^{3}$, then at $10^{3.25}$ etc. until $10^{6.75}$), resulting in an increasing number of revisions available for training. We repeat this exercise 10 times (mean $\pm$ stdv. shown in the Figure). For comparison, we depict the performance of the models trained on the all-topic datasets in all other languages (shown in blue). Note that in these cases there is no sampling, and for every language, all available data is used. We find that the performance of the English-language model is comparable to that of the other languages when there is little data available (up until about $10^{5}$ revisions) and is about 10% higher than the other languages when almost all data is used for its training. At least in the English-language case, we find that for stable model performance around 10,000 revisions are necessary. At the higher range of data, we find that the classifier’s performance has not levelled off, suggesting that adding more articles (possibly other topics) may improve the performance further.
Turning to editorial behaviour, the introduction of the perennial sources list provides a “natural experiment” to measure a possible editorial behaviour change in response to these additional guidelines. We only examine the impact of this change in the English-language datasets, as it is not clear whether editors from other language editions of Wikipedia used the English perennial sources list. Thus, we created a version of our English-language dataset using only revisions up to 2018-06-30 (i.e., before these guidelines were introduced). To control for the dataset size (which as shown above, may affect model performance), we perform again sub-sampling similar to what we did with the full English-language dataset. The black dash-dotted line and the grey area in Figure 4 show the model performance based on this limited data. We observe its performance to be notably lower than with more recent data (red line and area), indicating that the editorial behaviour before the introduction of the perennial sources list was less predictive of reliable/unreliable domain classes than afterwards. In this case, the performance is similar to that of other languages, showing that that editor behaviour allows extracting meaningful and predictive signals for source reliability even when the editors did not have an official guideline for source reliability. Furthermore, these findings support previous work which shows that the quality of the Wikipedia references improved after the introduction of the perennial sources (Baigutanova et al., 2023a).
4.3. RQ3: Model Adaptation across Topics & Languages
The multilingual nature of Wikipedia presents an opportunity to supplement information available in lower-resource languages with that in the larger ones. Considering our datasets, two directions of adaptation are possible: across languages and across topics. We consider both scenarios in Figure 5, which shows the performance of the models adapted in these two ways: in a cross-language setting (trained on a language and afterwards tested on a different language from the same resourcefulness class, in the same topic) and a cross-topic setting (trained on a topic, tested on a different topic, in the same language). We find that, in both cases, the performance of the adapted models decreases compared to the native performance. The decrease in performance is about the same in cross-language and cross-topic settings. This suggests that the editor’s behaviour around references may be non-identical in different topics, as well as languages. However, again, we see the importance of resourcefulness on the performance: the models adapted across high-resource languages perform better than those across mid and low-resource ones.
Figure 5. Model performance in two adaptation scenarios: cross-language setting (trained on a language, tested on different language from the same class, in the same topic, in red) and a cross-topic setting (trained on a topic, tested on a different topic, in the same language, in green). Results aggregated by language resourcefulness and compared to same-topic and language native models (in blue).
[…]
[¡short description¿]¡long description¿
Finally, we consider all topics together and analyse the performance of our model when trained with different training strategies and tested on each language. In particular, we test whether models trained on higher-resource data would be helpful to those in lower-resource settings. In Figure 6 we aggregate the corresponding results by languages of similar resourcefulness: the top panel shows how for many languages the different models perform significantly better, same as, or worse than a random model (according to Mann-Whitney tests with Bonferroni correction), the center panel shows the same, but compared to the native model, and the bottom panel shows the distribution of F1 macro scores. We find that F1 scores are generally lower when training a model on the English dataset and applying it to low-resource languages, confirming that cross-language adaptation can be problematic, even if a high-resource language such as English is used. However, when training on all high-resource languages or all languages together, the performance for low-resource languages significantly increases. This suggests that including a plurality of languages in training data may capture a variety of user behaviours that are transferable to lower-resource languages.
Nonetheless, each language-specific dataset has its own biases connected to the amount of data available, temporal peculiarities of the articles and their edits, and the volume of editorial activity. To alleviate potential difficulties in adapting feature values between languages, we normalize each dataset using quantile normalization (converting feature values to ranks). This approach further improves the performance of the all-language model adaptation, even on average performing better than the native models. Finally, since the size of the dataset increases with the combination of topics, we train the classifier with the maximum tree depth of 3 instead of 1 but do not find a substantial increase in performance.
Figure 6. Combined Model performances for different training datasets and strategies. Top panel: % of models that perform better, same as, or worse than a random classifier; Center: same, but compared to the native model; Bottom: violin plots of the corresponding distributions of the F1 macro performances.
[…]
[¡short description¿]¡long description¿
4.4. Qualitative Insights on Source Usage
The above quantitative evaluation of the proposed model abstracts the rich insight that its output provides. When examining the domains that attain the highest reliability score by the classifier, we find interesting signals of the editorial treatments of different sources, in each topical sphere (here, we focus on English data). For instance, among the domains predicted (incorrectly) as reliable in the Climate Change domain is YouTube (which appeared in 644 articles in that dataset) with a positive (reliable) class probability of $P_{r}=0.64$, as well as Internet Movie Database (IMDB) (appearing in 56 articles, $P_{r}=0.56$). This points to the importance of popular or mainstream culture in this topic. When considering the History dataset, we find the classifier incorrectly labels as reliable domains pointing to sources that the perennial sources label as having partisan bias, such as the Jewish Virtual Library ($P_{r}=0.81$), Sixth Tone ($P_{r}=0.74$), and Daily Sabah ($P_{r}=0.69$).
On the other hand, one can use the output of this classifier to find sources that different communities of editors consider inappropriate for their topical domain. For instance, in the Biology dataset, Science-Based Medicine, which is considered generally reliable and has “a credible editorial board” according to the perennial source list, is treated by the editors more similarly to unreliable domains ($P_{r}=0.20$). Similarly, the Huffington Post, rated as “fairly reliable for factual reporting on non-political topics”, scores low in the COVID dataset ($P_{r}=0.07$), possibly due to the politically charged and serious nature of the rhetoric around the pandemic.
4.5. Alternative Sources of Reliability
One limiting factor of this study was the availability of labels for the quality of references. We consider an alternative: Media Bias / Fact Check (MBFC), an “independent website” that provides ratings of “the bias, factual accuracy, and credibility of media sources” [^8] https://mediabiasfactcheck.com/about/, which has been widely used in misinformation research and by other indexes (Barret, 2021; Chowdhury et al., 2020; Ye and Skiena, 2019). We use this credibility metric and consider very high and high as the positive class, and low and very low as negative. With the same features and model as above and LOO cross-validation, we gauge its performance on MBFC. Although the class-average F1 macro is 0.61, the classifier performs especially poorly on the negative class, with a precision of 0.32 and recall of 0.55, suggesting that the features do not capture this metric. Indeed, there may be several reasons why such resources are not suitable for the analysis of Wikipedia references. First, there is little overlap between the sources noted in MBFC and in the English Wikipedia perennial sources list (out of 1156 perennial sources, 272 are in MBFC). Second, the MBFC credibility metric is weakly aligned with the perennial source reliability one. Third, MBFC itself is listed as generally unreliable by the English Wikipedia perennial sources list.
5. Discussion
In this study, we have proposed and assessed editing behaviour-based features for modeling the reliability of web domains used as references on Wikipedia. Editing behaviour is not the only source of useful information for language-agnostic modeling of source reliability. Links between articles (already used for topical clustering of Wikipedia articles (Johnson et al., 2021)) may provide semantically meaningful context for references, whereas users’ reading sessions (used for entity linking (Gerlach et al., 2021)) may give controversiality signals from the perspective of the audience. Future work will also take a page from the social media literature and build editor-resource networks (Shu et al., 2019) that would capture structural peculiarities of topics, editors, and resources associated with low-quality content.
Apart from finding that the proposed features capture editorial behaviour within a dataset spanning one topic and language, we also observe that they retain useful information during the model adaptation across languages. Although English has been the lingua franca in NLP, having the best-developed tools and resources, we show that adaptation from the English-language dataset to other languages is largely inferior to using a model that combines information from many languages. This improvement is especially noticeable in performance on low-resource languages. These findings are aligned with limitations posed by Das et al. (2024) suggesting that, although the features are language-agnostic, limiting data selection to one language may limit the breadth of the captured behaviours and their applicability to new settings. As multilingual embedding models have gained popularity (Zhang et al., 2023a), it remains to be seen whether the latest developments in NLP will benefit the low-resource languages (Zhang et al., 2023b), and what the impact will be on Wikipedia specifically. [^9] https://en.wikipedia.org/wiki/Wikipedia:Why_you_shouldn%27t_write_articles_with_ChatGPT,_according_to_ChatGPT The need to monitor possible changes in the editorial standards and accompanying quality of new references calls for further development of features reflecting these behaviors, such as those described in the present study.
As ground truth, we have used the list of perennial sources (Wikipedia contributors, 2024) from the English Wikipedia, which is part of Wikipedia’s efforts to provide clear editorial instructions to its community. We compare this resource with another popular source of reliability measure, the MBFC (Zandt, 2024), and find significant differences between them, pointing to the necessity of Wikipedia-specific models like ours. Previous work comparing the perennial sources lists in different languages showed disagreement on whether some domains are indeed reliable (Baigutanova et al., 2023b). Therefore, the system we propose could be used to expand lists of perennial sources and lists derived from them [^10] https://science.feedback.org/consensus-credibility-scores-comprehensive-dataset-web-domains-credibility/, or even to assess cultural biases in Wikipedia sources (Zheng et al., 2023; Yang and Colavizza, 2024).
Limitations
The present study has multiple limitations. Using language-agnostic approaches may be seen as a limitation, as it does not take advantage of the predictive power of language-dependent features. However, despite the significant advances with multilingual NLP resources such as mBERT (Devlin et al., 2018), the motivation for our model was precisely to be ready for use in the 300+ language editions of Wikipedia, especially those that are less privileged and under-resourced.
Second, while the dataset contains thousands of articles and cumulatively millions of revisions, it spans only a fraction of Wikipedia. Despite choosing topical foci such that different reference quality and time scales, they may not capture the entire diversity of subject matter of this massive collaborative effort. For instance, differentiating topics by whether the topic is currently controversial may reveal shifts in editorial handling of sources.
Third, the selection of the English perennial sources list as the gold standard disadvantaged other languages. Unfortunately, only a few other languages have a similar resource and they are significantly less extensive and conclusive (Baigutanova et al., 2023b). Given the global role played historically by the English edition of Wikipedia (Hale, 2014), our selection provided the best choice to reach the maximum potential coverage across other language editions. In the future, we aim to use the scores computed by the adapted models as initial input for future collaboration with the editors of these language editions to create language-specific source quality guidelines.
Finally, we show that the method is sensitive to the amount of data available for training (and testing), possibly resulting in insignificant results in some low-resource language settings. Further research will be needed to extend the amount of available data in these contexts.
Ethical Considerations
This study, and online content moderation in general, have important ethical aspects that should be kept in mind both by the platforms and the misinformation researchers.
First, this study has a direct impact on the access equity to encyclopedic resources in low-resource languages. Currently, the lack of resources such as the perennial source list for these versions of Wikipedia limits the extent to which our model is able to reflect the standards of that linguistic community. The domain scores provided by the models adapted from other (higher resource) languages that we propose here could be a starting point for bootstrapping the editorial discussion of the quality of sources used in low-resource languages.
Second, the process of determining the quality of sources bears scrutiny. In the news domain, the quality of information is often determined by professional fact-checkers. However, on Wikipedia, the editors from the community define the status of the sources via a deliberation process, guided by the community standards. Any potential biases in the judgment of these editors or failings in the deliberative process would be reflected in the quality of these labels (for instance, according to Kharazian et al. (2023), some Wikipedia editions were dominated for a time by a “small group of administrators who introduced far-right bias and outright disinformation”). Generally, in 2014 Shaw and Hill (2014) find that the Wikipedia contributors self-organize into oligarchic structures, though in 2021 Rijshouwer et al. (2023) find “strong counter-tendencies” to editorial power concentration. Our proposed method may reveal community behaviors which are not necessarily explicitly stated, but which shape the platform’s use of outside sources, potentially aiding the community to self-monitor the use of sources.
We believe that the code and resources we make available to the community [^11] https://github.com/JacopoDignazi/Wiscom (compiled in accordance with FAIR (Findable, Accessible, Interoperable, and Reusable principles (Wilkinson et al., 2016)) will be useful to assist future research efforts and to further evaluate and extend the features and modeling we proposed.