### v3

# Language-Agnostic Modeling of Source Reliability on Wikipedia

Abstract.
Over the last few years, verifying the credibility of information sources has become a fundamental need to combat disinformation. Here, we present a language-agnostic model designed to assess the reliability of web domains as sources in references across multiple language editions of Wikipedia. Utilizing editing activity data, the model evaluates domain reliability within different articles of varying controversiality, such as Climate Change, COVID-19, History, Media, and Biology topics. Crafting features that express domain usage across articles, the model effectively predicts domain reliability, achieving an F1 Macro score of approximately 0.80 for English and other high-resource languages. For mid-resource languages, we achieve 0.65, while the performance of low-resource languages varies. In all cases, the time the domain remains present in the articles (which we dub as permanence) is one of the most predictive features. We highlight the challenge of maintaining consistent model performance across languages of varying resource levels and demonstrate that adapting models from higher-resource languages can improve performance. We believe these findings can assist Wikipedia editors in their ongoing efforts to verify citations and may offer useful insights for other user-generated content communities.
Wikipedia, knowledge equity, knowledge integrity, information reliability
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXccs: Information systems Wikisccs: Computing methodologies Machine learning algorithms
1. Introduction
The proliferation of fake news in recent years is deteriorating the reliability and trustworthiness of the online information ecosystem (Del Vicario et al., 2016; Vosoughi et al., 2018). To mitigate their spread and impact in society, researchers in web and data mining are devoting a great deal of effort to generating resources to detect untrue content on social media platforms (Su et al., 2023; Xu et al., 2022; Ashkinaze et al., 2024; Xiao et al., 2024; Guo et al., 2023b; Xiao et al., 2023; Wu et al., 2023).
In popular web platforms like Wikipedia [^1] Wikipedia is among the top visited websites worldwide: https://www.semrush.com/trending-websites/global/all., disinformation is also a main threat to knowledge integrity (Aragón and Sáez-Trumper, 2021). Wikipedia’s conception of knowledge as a service requires content to be verifiable to preserve its reliability and integrity [^2] https://research.wikimedia.org/knowledge-integrity.html. Thus, the platform’s policy on verifiability implies that readers should be able to check that all the information in Wikipedia articles comes from reliable sources [^3] https://en.wikipedia.org/wiki/Wikipedia:Verifiability.. While a classical patrolling technique in this platform is to detect and remove statements that violate basic core content policies (Kumar et al., 2016), some Wikipedia editors have proven more effective in tackling disinformation by first identifying unreliable sources (Cohen, 2021). Recent research revealed the positive impact of the community-curated list of perennial sources in English Wikipedia (Baigutanova et al., 2023a). However, that list is limited or even non-existent in other language editions (Baigutanova et al., 2023b). A great deal of manual effort would be necessary to create such community-curated resources in all of the languages on Wikipedia.
To address this challenge, we propose a language-agnostic approach to reveal source reliability that leverages the implicit signals in the edit activity data from multiple Wikipedia language editions. We aim to support Wikipedia editors, including those in medium and small projects that often miss advanced tools, in identifying sources with patterns associated with low reliability, and in monitoring their prevalence across language editions. To meet this goal, we address the following research questions:
  • RQ1. What are the most predictive language-agnostic features when modeling source reliability in Wikipedia, considering the Climate Change topic in English as a case study?
  • RQ2. To what extent can model performance be related to the topic and size of a language edition of Wikipedia?
  • RQ3. Is it possible to adapt these models across topics and/or languages?
Our methods to answer these questions are inspired by existing language-agnostic approaches to measure the controversiality of linked elements in a given Wikipedia article (Borra et al., 2015). Results with datasets from different topics and language editions suggest a promising scenario. Given the several limitations outlined in the following sections, we pay attention to model performance not only for the largest editions but also in mid- and low-resource settings, where adaptation of the models may provide the most benefit. To our knowledge, this is the first work modeling source reliability on Wikipedia using only language-agnostic features.
2. Related Work
This work approaches the challenge of measuring source reliability on Wikipedia through the lens of controversiality metrics. This approach builds on the methodology of the Contropedia project (Borra et al., 2015), aimed at identifying and visualizing controversial links between Wikipedia articles, i.e., links that point to other articles in the platform and whose collaborative edit history shows signs of dispute between editors before reaching consensus. This is not the only effort in this space. Systems like Edit-History Vis (Guo et al., 2023a) have also included interactive visualizations to illustrate the path to consensus for a Wikipedia article. Cultural differences in the controversiality of specific topics have been shown to exist (e.g., (Rogers et al., 2012; Fu et al., 2023)). Unlike these previous approaches primarily focusing on wikilinks within article text, our work shifts attention to citations, aiming to explore the reliability of sources in Wikipedia.
Previous research has examined assessed the verifiability of Wikipedia content (Chen and Roth, 2012; Benjakob et al., 2022; Halitaj and Zubiaga, 2024, 2025) and the reliability of the sources (Lewoniewski et al., 2020; Yang and Colavizza, 2024).  Baigutanova et al. (2023a) addressed both goals by analyzing reference quality on English Wikipedia over time using two key metrics: the proportion of sentences lacking necessary citations, as identified by machine learning models (Redi et al., 2019; Chou et al., 2020), and the proportion of citations linking to non-authoritative sources. For the latter, non-authoritative sources were determined using the perennial source list (Wikipedia contributors, 2024), a curated list of sources whose reliability and use on Wikipedia are frequently discussed. The study found the share of references to non-authoritative sources has remained below 1%, with a notable decline following the introduction of the list in 2018.
Although the perennial source list proved effective in decreasing the presence of non-authoritative sources on English Wikipedia, later research revealed significant limitations when applying this concept to other language editions (Baigutanova et al., 2023b). A few other language communities have developed their own versions of the perennial source list, but such cases are rare among the hundreds of Wikipedia editions, and these lists typically contain only limited data. Furthermore, sources flagged as unreliable in one language edition are often still used in others, highlighting the variability in how source reliability is perceived across cultural and linguistic contexts. These discrepancies pose a major challenge to the development of universally applicable models for assessing source reliability across Wikipedia language editions.
The multilingual nature of Wikipedia calls for specific modeling techniques (Johnson and Lescak, 2022). Models such as multilingual BERT have been used to detect vandalism (Trokhymovych and Saez-Trumper, 2021) or political bias (Swati et al., 2023), while other works have built rankings of relative quality in multiple language versions of the platform (Lewoniewski et al., 2020, 2019). To build a multilingual model, such as one recently proposed for entity insertion (Feith et al., 2024), one needs a large dataset of training data in all the languages of interest. For some other tasks, and due to the limitations found in languages with smaller Wikipedia datasets, language-agnostic approaches. These approaches have proven effective across a wide range of tasks, e.g., the assessment of article content quality (Das et al., 2024), reliability (Wong et al., 2021), automatic entity-linking (Gerlach et al., 2021), orphan article detection (Arora et al., 2024), and topic modeling (Piccardi and West, 2021) and classification (Johnson et al., 2021).
Language-agnostic models using features extracted from the network structure or user behavior have also been used in the more general context of detecting misinformation or low-quality information in social media. In this line, Shu et al. (2019) constructed embeddings of users and news and used diffusion models to trace the paths through which fake news propagates, while Zhao et al. (2021) extracted features from the user interaction network to help identify misinformation in online health communities. As shown later, user-behavior related language-agnostic signals are useful for modeling source reliability, and can be adapted across languages to improve performance in low-resource settings.
3. Data and Methods
3.1. Perennial Sources
We use the English Wikipedia perennial source list (Wikipedia contributors, 2024) as ground truth for modeling source reliability. As only a few other languages have perennial source lists, in this work we focus on the list from English Wikipedia, as it is the most complete one, and concerns the largest language Wikipedia. We discuss this lack of ground truth data in the Limitations section below. It was collected on March 30, 2023. Although the list exists for 12 other languages, we only use the English version as it is by far the most extensive (see (Baigutanova et al., 2023b) for details on coverage). In English, it consists of a collection of web domains in five categories: blacklisted (569), deprecated (125), generally unreliable (178), no consensus (137), and generally reliable (184). To build a binary classifier, we use the generally reliable domains as positive labels, merge the domains blacklisted, deprecated, and generally unreliable into a single unreliable category (i.e., the negative label), and ignore the no consensus domains.
Table 1. Statistics of the topic-language datasets used in the study grouped by high, mid, and low resource languages. All numbers are averages (apart from the number of languages). Perennial Domains (P. Domains) refers to the number of domains for the perennial source list in English found in the different datasets. See Table A1 in Appendix B for the corresponding median values.
Topic Langs Articles Revs Revs/Ar URLs URLs/Ar Domains P. Domains
High Climate Change 7 $1378$ 372 133372\,133 $269$ 58 67158\,671 $49$ 15 36315\,363 $203$
COVID-19 7 $898$ 233 905233\,905 $260$ 60 16860\,168 $89$ $8289$ $206$
Biology Sample 7 13 60713\,607 586 171586\,171 $43$ 59 76359\,763 66 11 71311\,713 $148$
History Sample 7 $5710$ 624 428624\,428 $109$ 52 70252\,702 $12$ 12 12112\,121 $166$
Media Sample 7 $5577$ 819 541819\,541 $146$ 108 583108\,583 $22$ 21 29021\,290 $259$
Mid Climate Change 36 $395$ 28 36428\,364 $71$ $6224$ 1919 $2599$ $94$
COVID-19 37 $272$ 18 52318\,523 $68$ $8984$ $39$ $1788$ $102$
Biology 37 $4153$ 83 59983\,599 2020 $8691$ 33 $2022$ $61$
History 37 $1680$ 83 25883\,258 $49$ $6274$ 77 $2054$ $67$
Media 37 $1018$ 46 95746\,957 $46$ $7932$ $10$ $2416$ $123$
Low Climate Change 56 $115$ $3430$ $29$ $1079$ $12$ $570$ $35$
COVID-19 75 $32$ $916$ $28$ $568$ $22$ $242$ $32$
Biology 41 $593$ 21 09621\,096 $35$ $1159$ 44 $460$ $16$
History 55 $511$ 13 50113\,501 $26$ $952$ 66 $406$ $22$
Media 70 $159$ $2262$ $14$ $528$ 77 $264$ $30$
3.2. Language Groups
As we aim to test the proposed model on a variety of languages, we collect data on all $326$ available language editions of Wikipedia. We then discard all languages that have fewer than two reliable and unreliable sources according to the English perennial source list. Among the remaining ones, we distinguish between low, middle, and high-resource languages by examining the number of “active users”, defined as “registered users who have made at least one edit in the last $30$ days.” [^4] https://meta.wikimedia.org/wiki/List_of_Wikipedias (accessed on 01-03-2024) We consider the top 5% of the languages having the most active users as high-resource, the subsequent 25% as mid-resource, and the remaining 70% as low-resource languages.
3.3. Article Collection
3.3.1. Topic Selection
In this study, we have focused on articles from the following five topic datasets: Climate Change, COVID-19, Biology, History, and Media. Climate Change and COVID-19 were selected as they are both controversial, but for different time durations. COVID-19 appeared at the end of 2019, while Climate Change has been a controversial topic for decades (Korte et al., 2023). Both have a Wikiproject community of editors  [^5] https://en.wikipedia.org/wiki/Wikipedia:WikiProject who label articles as relevant to the topic with special guidelines. For comparison, we also selected three topics that are not as controversial, but which varied in terms of reference quality based on the “reference risk” metric. This metric was introduced by Baigutanova et al. (2023a) and estimates the likelihood of articles citing unreliable sources. Specifically, reference risk is calculated as the proportion of citations in an article that point to domains categorized as generally unreliable, deprecated, or blacklisted in the English Wikipedia perennial source list. Higher reference risk indicates a greater potential threat to verifiability and information quality.
This stratification allowed us to evaluate our model’s performance across topics with varying levels of citation quality and editorial scrutiny.
3.3.2. Data Download
For the first two topics, we use the articles listed by the corresponding topic-based WikiProjects. For the latter three, we use the topics identified using the Wikipedia Objective Revision Evaluation Service (ORES) machine learning system (Halfaker and Geiger, 2020). We use this method rather than later approaches (Johnson et al., 2021), since it is the one employed in Baigutanova et al. (2023a), which inspires our comparison among topics. To keep the sizes of datasets comparable and manageable, we sample the articles identified by ORES to be in Biology and History topics by 10% (each), and in the Media topic by 2%. These two approaches allow us to find articles in the English Wikipedia for each of the five topics, resulting in a total of 84 02784\,027 articles. To find the versions of these articles in other languages, we use the MediaWiki API, which results in another 335 809335\,809 of articles in all of the 326 languages. We collect the articles and their complete edit history using the MediaWiki API in between December 2023 and February 2024.
3.3.3. Extraction of Source Domains
Citations may appear in the body of the article text and as references at the bottom of the article, each of which has a standard template. Each revision that involves a citation contains the author of the revision, position within the document of the citation, text of the citation, and meta information, such as whether it is a DOI, URL, ISBN, ISSN, etc. or a raw URL. We standardize these revisions to create a list of “source edits” containing the metadata about each insertion, deletion, or edit that involves a source. We resolve redirects, apply rules to clean and standardize the URLs (including DOI references redirecting to a resource), and merge revisions when the same user edits an article in succession (regardless of the time). Finally, we extract the domains from the URLs, obtaining 376 566376\,566 unique source domains.
3.3.4. Dataset Statistics
The summary statistics about the topic-language datasets (or simply “datasets” throughout this paper), for the five topics grouped into low-, mid-, and high-resource languages can be seen in Table 1. The medians of these values can be found in Appendix B, while more in-depth statistics for each individual language are in Appendix C. Recall that we only use datasets that have passed the threshold of having at least 2 reliable and 2 unreliable perennial sources.
The smallest topic in terms of the number of articles is COVID-19, while the highest numbers of revisions per article are found around Climate Change and COVID-19 in high-resource languages. On the contrary, biology has the largest number of articles, but these are revised on average the fewest number of times in high- and mid-resource languages. High-resource languages have much better coverage of the English perennial sources, with an average of 150 to 260 sources per topic. These values decrease to an average of 15 to 35 sources for low-resource languages, making the training and evaluation more challenging.
3.4. Feature Definition
Unlike previous studies measuring Wikipedia edit quality that did not consider certain types of edits and reverts (Trokhymovych et al., 2023), we propose to focus on the entire edit history concerning each domain. Using edit metadata, we define 52 features capturing the popularity of the domain, the “permanence” of the domain across edits, and the number and type of users involved in edits adding or removing that domain. As the features are computed for each domain within each dataset separately (recall, by “dataset” we mean a collection of articles on a certain topic, in a certain language) and the datasets are of varying sizes and editorial frequencies, we explore different ways of normalizing these statistics w.r.t. the first appearance of the domain, age of the dataset, and in terms of time duration or number of revisions. Below, we give a short overview of the features we have used. A more detailed description can be found in Appendix A.
3.4.1. Popularity features
  • $N_{\mathrm{articles}}$, $\overline{N_{articles}}$: Number of articles a domain has appeared in (sensitive to the size of the dataset), and its normalized version (by the number of articles in the dataset).
  • $CurrN_{articles}$, $\overline{CurrN_{articles}}$: Number of articles a domain is used in at collection time, and normalized by the total number of articles in the dataset.
3.4.2. Permanence features
  • $\Sigma Perm_{d}$, $\Sigma Perm_{r}$: Permanence, how long a domain has been used in an article (not necessarily consecutively), summed across all articles and measured in days (subscript $d$) or by this number of revisions (subscript $r$).
  • $\Sigma CurrPerm_{d}$, $\Sigma CurrPerm_{r}$: Same as above, but considering only the articles where the domain is currently used.
  • $\overline{\Sigma Perm_{d}}$, $\overline{\Sigma Perm_{r}}$: Sum of all permanences for a domain in all articles, normalized by the sum of the ages of all articles (measured by the number of days or the number of revisions that an article existed).
  • $\overline{\Sigma CurrPerm_{d}}$, $\overline{\Sigma CurrPerm_{r}}$: Same as above, but considering only the articles where the domain is currently used.
  • $\langle Perm_{d}\rangle$, $\langle Perm_{r}\rangle$: The average permanence of a domain (over all articles where it was used).
  • $\langle SelfPerm_{d}\rangle$, $\langle SelfPerm_{r}\rangle$: Self-permanence: permanence divided by the number of days (or revisions) since the first time the domains were added to the article, then averaged article-wise.
  • $\Sigma age_{d}$, $\Sigma age_{r}$: The sum of ages of the domain over all the articles in a dataset, where the age of a domain in an article is the number of days or revisions since a URL from that domain has been first added to the article.
  • $\langle age_{d}\rangle$, $\langle age_{r}\rangle$: The average age of a domain over all the articles it appears at least once.
3.4.3. User-based features
  • $U_{add}$, $U_{rem}$: Number of users that have added or removed a domain.
  • $R_{add}$, $R_{rem}$: Same as above, but only counting registered users.
  • $\overline{U_{add}}$, $\overline{U_{rem}}$, $\overline{R_{add}}$, $\overline{R_{rem}}$: The above four features, normalized by the total number of unique users in the dataset.
  • $\langle U_{add}\rangle$, $\langle U_{rem}\rangle$, $\langle R_{add}\rangle$, $\langle R_{rem}\rangle$: The above four features, measured as an average per article.
  • $Ratio(R_{add}/U_{add})$, $Ratio(R_{rem}/U_{rem})$ The proportion of registered vs. all users that ever added or removed a domain.
  • $Proba(R_{add})$, $Proba(R_{rem})$ Probability that, when a domain is added or removed, this is done by the registered users. Contrary to the ratio, here we take into account the number of revisions per user.
  • For all user-based features listed above, we compute a version wherein instead of counting any instance of adding a domain, we only count instances of adding it for the first time on an article (“starting”), and similarly, counting only the instances that the domain was removed for the last time (“ending”). For example, $Proba(R_{add})$, and $Proba(R_{rem})$ would have a corresponding version $Proba(R_{start})$ and $Proba(R_{end})$ (similarly for the related features).
We remark that, to aggregate the various statistics over articles for a domain, we have considered both the average and the median and found the best performance with the average, which we use in the models in the following sections. [^6] We have also experimented with versions of the above features that use the aggregation over all URLs in each dataset, instead of domains. However, the resulting features were performing very similarly to those aggregated over domains, and we excluded these from the analysis.
3.5. Model Training and Evaluation
Using the above features, we train an XGBoost [^7] We use Python package XGBoost V1.6.2 https://pypi.org/project/xgboost/1.6.2/. model to classify the domains as reliable or unreliable. To account for class imbalance, we use weights (e.g., the weight for the positive class is the fraction of negative/positive). To limit overfitting, we set the maximum depth of the classifier to 1 (we have tested the classifier with depths up to 5, but found no increase in leave-one-out (LOO) validation performance, and the learning rate to $\eta=0.1$ (default is 0.3).
When evaluating in the “native” condition, wherein the training and target data are from the same dataset, we employ LOO cross-validation. In the “cross-language adaptation” condition, wherein the model is trained on one language and tested on another (keeping the topic the same), we simply compute the evaluation metrics on the test language. Similarly, for cross-topic adaptation, we keep the language consistent and test on a target topic. In the “mixed” condition, wherein the model is trained on a set of languages and tested on one of these languages, we again employ LOO cross-validation where a domain of test language is left out at a time.
To compute a confidence interval for a metric, we employ bootstrapping wherein we resample with replacement the output of the classifier and recompute the metrics $n=100$ times. The main metric we compute is the F1 macro score, which is the unweighted average of the F1 scores of the negative and positive classes.
4. Results
Table 2. Leave-one-out validation F1 macro, and precision and recall metrics for both classes for the model trained and tested in the English Wikipedia on each topic, and for all topics combined (standard deviations specified with $\pm$). Class balance in terms of the number of reliable (rel) and unreliable (unr) domains.
Topic F1 Macro Precision (rel) Recall (rel) Precision (unr) Recall (unr) # rel # unr
Climate Change $0.81\pm 0.02$ $0.80\pm 0.03$ $0.80\pm 0.04$ $0.83\pm 0.04$ $0.83\pm 0.03$ $129$ 152
COVID-19 $0.88\pm 0.02$ $0.87\pm 0.03$ $0.91\pm 0.03$ $0.89\pm 0.03$ $0.85\pm 0.03$ $136$ 125
Biology $0.80\pm 0.03$ $0.81\pm 0.03$ $0.81\pm 0.03$ $0.80\pm 0.03$ $0.80\pm 0.03$ $124$ 120
History $0.75\pm 0.03$ $0.70\pm 0.04$ $0.83\pm 0.04$ $0.82\pm 0.03$ $0.69\pm 0.05$ $120$ 139
Media $0.83\pm 0.02$ $0.81\pm 0.03$ $0.82\pm 0.03$ $0.85\pm 0.03$ $0.84\pm 0.03$ $146$ 176
All topics $0.83\pm 0.02$ $0.76\pm 0.04$ $0.85\pm 0.03$ $0.89\pm 0.02$ $0.82\pm 0.03$ $159$ 234
Figure 1. Beeswarm analysis of an XGBoost model trained on Climate Change English dataset (class encoding is 1: reliable and 0: unreliable) and two example classification explanations (score shown is the log-odds ratio). : ¡short description¿
4.1. RQ1: Case study: Climate Change in English
We begin by studying the behavior of our model for the Wikipedia articles in English around the Climate Change topic. The model performance is shown in the first row of Table 2, which lists the F1 Macro, precision and recall for each class. We find that, out of the unreliable domains, the classifier correctly identifies as unreliable 83% of them (recall), and out of the domains the classifier guessed as unreliable, 83% are truly unreliable (precision). Both precision and recall for the reliable class are slightly worse at 80%. Note that the classes are fairly balanced in this setting, with 46% of the dataset having the reliable label.
Figure LABEL:fig:RQ2_beeswarm shows the “bee swarm” plot of SHAP (SHapley Additive exPlanations) values that illustrate the importance of each feature in the classifier (Lundberg and Lee, 2017). The top 6 most predictive features (and an aggregation of the rest) are shown. The SHAP values for the top feature $Proba(R_{end})$ – the probability that, when a domain is removed, it is done by a registered editor – reveal that high values of this feature are associated with a negative impact on the domain reliability prediction. In other words, when a registered editor removes a domain, it is more likely to be unreliable. The opposite can be seen for the $\langle SelfPerm_{d}\rangle$ –average proportion of days the domain remains cited since it was first added– and $\langle Perm_{r}\rangle$ –the average number of revisions that the domain remains on the article–: higher values of these features are associated with more reliable sources. $\Sigma CurrPerm_{r}$ –the aggregated number of revisions through which the domain has existed across all articles– also correlates positively with the reliable category, as does as well the sum of the domain ages across articles, $\Sigma age_{r}$. Finally, “younger” domains – those that have been added to an article more recently, $\langle age_{d}\rangle$ – are more likely to be reliable. This may be a sign of the increased focus of the platform on the quality of sources (Baigutanova et al., 2023a).
In Figures LABEL:fig:RQ1_twitter and LABEL:fig:RQ1_Researchgate we illustrate the performance of the model for two domains of special interest: twitter.com, correctly identified as unreliable, and researchgate.net, incorrectly identified as reliable. In the case of twitter.com, despite being added/present frequently on articles (high $\Sigma CurrPerm_{r}$), it is soon removed by registered editors ($Proba(R_{end})$ and also SelfPermanence, $\langle SelfPerm_{d}\rangle$ is low). For researchgate.net, on the other hand, model and ground truth disagree: most of the permanence features indicate that it is used as a reliable source, while the domain is considered unreliable in the perennial source list because of being a self-publishing platform lacking editorial oversight or peer review. However, as our data points out, Wikipedia editors may still use it pragmatically, for example, to link to freely accessible versions of peer-reviewed papers that are otherwise behind paywalls. This example underscores the potential of our method to reveal discrepancies between Wikipedia’s formal sourcing guidelines and how editors apply them in practice, highlighting edge cases that may warrant further discussion or refinement in policy.
4.2. RQ2: Topic and Language Generalizability
We continue by evaluating the performance of our language-agnostic modeling approach in other topical and linguistic settings. Table 2 shows the F1 macro and class-specific precision and recall for these topics, in the English Wikipedia. We find that, overall, the performance is consistent, with the lowest performance observed for the History dataset, with a recall of 0.69 for the unreliable class. The best performance is achieved on the COVID-19 dataset, possibly due to the ample data on the edits and the more recent nature of the corresponding articles. We find that the amount of edited data is important for the classifier’s performance, as we show later in this section. Finally, we combine all English-language datasets into one and perform training and LOO validation; the performance of this model is shown in Table 2 under “all topics”. Such a combined model performs on par with other topical models, with especially high precision for the unreliable class (0.89), suggesting that combining knowledge about different topics does not overall degrade the performance of the classifier.
Next, we explore the applicability of the proposed model to other languages. In these experiments, we first keep the topic constant – Climate Change – while training and evaluating models in different languages. Figure 2 shows the LOO F1 macro scores for select languages form the high-, mid-, and low-resource groups, in comparison to the performance of a random baseline model. Specifically, we show the average performance $\pm$ one standard deviation as shaded areas. To estimate the performance of a random model, we assign a random binary classification for the domains and perform bootstrapping ($n=100$) similar to the evaluation of all other models, as discussed in Section 3.5. The performance is highest for high-resource languages and degrades for mid- and low-resource ones. The standard deviations of the scores also increase with lower resource languages (both for our model and the baselines), pointing to an increased lack of information for computing the score.
Additionally, we show aggregate statistics of the model performance on the languages in each category in Figure 3. In particular, we are interested in whether our model outperforms the random classifier in terms of F1 macro score (top panel of the figure): we perform this comparison using Mann-Whitney scores and adjust the significance level of 0.05 using the Bonferroni correction for multiple comparisons. For instance, in the Climate Change topic, the model outperforms the random baseline in 100% of the high-resource languages, but only in 44% of low-resource languages (in an additional 12% the difference is not statistically significant). The performance on low-resource languages tends to vary widely for different topics. The brown bars show the models for which articles for all topics within a language are used for training and testing. These models perform markedly better than topic-specific ones, echoing our findings for the English language.
Figure 2. Performances (mean $\pm$ stdv) of models trained and validated in languages other than English on the Climate Change datasets. The three batches of languages, separated by a vertical black line, show results on a sample of high (left), mid (center), and low (right) resource languages.: ¡short description¿
Figure 3. Native model performance per topic for different language types. Top panel: % of models that perform better than a random classifier. Bottom panel: Violin plots of corresponding distributions of F1 macro performances.: ¡short description¿
Figure 4. Average model F1 macro (line) and standard dev. (shaded area) versus the size of the training dataset in terms of the number of revisions. In English (red), all topics are combined, and the data is sampled at regular intervals. The same experiment is repeated, considering only revisions before the implementation of the perennial sources on 2018-07-01 (grey). The performance of models in all other languages (also all-topic) is shown in blue, without sub-sampling.: ¡short description¿
Effect of dataset size and introduction of perennial sources on model performance
The worse performance of our model in the lower-resource languages may be mainly (among others) due to two reasons: the datasets available for training the model are too small, or the behavior of the users of these languages towards domains is intrinsically difficult to model (their treatment of reliable and unreliable domains is difficult to distinguish). We design two experiments to address these two reasons, starting with the model’s sensitivity to the amount of training data.
The red line in Figure 4 shows the performance of the model trained on different dataset sizes. In particular, we consider the English (all-topic) dataset and sample the data at log-regular intervals (starting at $10^{3}$, then at $10^{3.25}$ etc., until $10^{6.75}$), resulting in an increasing number of revisions available for training. We repeat this exercise 10 times (mean $\pm$ stdv. shown in the Figure). For comparison, we depict the performance of the models trained on the all-topic datasets in all other languages (shown in blue). Note that in these cases, there is no sampling, and for every language, all available data is used. We find that the performance of the English-language model is comparable to that of the other languages when there is little data available (up until about $10^{5}$ revisions) and is about 10% higher than the other languages when almost all data is used for its training. At least in the English-language case, we find that for stable model performance around 10,000 revisions are necessary. At the higher range of data, we find that the classifier’s performance has not leveled off, suggesting that adding more articles (possibly other topics) may improve the performance further.
Turning to editing behavior, the introduction of the perennial sources list provides a “natural experiment” to measure a possible editing behavior change in response to these additional guidelines. We only examine the impact of this change in the English-language datasets, as it is not clear whether editors from other language editions of Wikipedia used the English perennial sources list. Thus, we created a version of our English-language dataset using only revisions up to 2018-06-30 (i.e., before these guidelines were introduced). To control for the dataset size (which, as shown above, may affect model performance), we again perform sub-sampling similar to what we did with the full English-language dataset. The black dash-dotted line and the grey area in Figure 4 show the model performance based on this limited data. We observe its performance to be notably lower than with more recent data (red line and area), indicating that the editing behavior before the introduction of the perennial sources list was less predictive of reliable/unreliable domain classes than afterwards. In this case, the performance is similar to that of other languages, showing that editor behavior allows extracting meaningful and predictive signals for source reliability even when the editors did not have an official guideline for source reliability. Furthermore, these findings support previous work, which shows that the quality of the Wikipedia references improved after the introduction of the perennial sources (Baigutanova et al., 2023a).
4.3. RQ3: Model Adaptation across Topics and Languages
The multilingual nature of Wikipedia presents an opportunity to supplement information available in lower-resource languages with that in the larger ones. Considering our datasets, two directions of adaptation are possible: across languages and across topics. We consider both scenarios in Figure 5, which shows the performance of the models adapted in these two ways: in a cross-language setting (trained on a language and afterwards tested on a different language from the same resourcefulness class, in the same topic) and a cross-topic setting (trained on a topic, tested on a different topic, in the same language). We find that, in both cases, the performance of the adapted models decreases compared to the native performance. The decrease in performance is about the same in cross-language and cross-topic settings. This suggests that the editor’s behavior around references may be non-identical in different topics, as well as languages. However, again, we see the importance of resourcefulness on the performance: the models adapted across high-resource languages perform better than those across mid and low-resource ones.
Figure 5. Model performance in two adaptation scenarios: cross-language setting (trained on a language, tested on a different language from the same class, in the same topic, in red) and a cross-topic setting (trained on a topic, tested on a different topic, in the same language, in green). Results aggregated by language resourcefulness and compared to same-topic and language native models (in blue). : ¡short description¿
Finally, we consider all topics together and analyze the performance of our model when trained with different training strategies and tested on each language. In particular, we test whether models trained on higher-resource data would be helpful to those in lower-resource settings. In Figure 6 we aggregate the corresponding results by languages of similar resourcefulness: the top panel shows how for many languages the different models perform significantly better, same as, or worse than a random model (according to Mann-Whitney tests with Bonferroni correction), the center panel shows the same, but compared to the native model, and the bottom panel shows the distribution of F1 macro scores. We find that F1 scores are generally lower when training a model on the English dataset and applying it to low-resource languages, confirming that cross-language adaptation can be problematic, even if a high-resource language such as English is used. However, when training on all high-resource languages or all languages together, the performance for low-resource languages significantly increases. This suggests that including a plurality of languages in training data may capture a variety of user behaviors that are transferable to lower-resource languages.
Nonetheless, each language-specific dataset has its own biases connected to the amount of data available, temporal peculiarities of the articles and their edits, and the volume of editing activity. To alleviate potential difficulties in adapting feature values between languages, we normalize each dataset using quantile normalization (converting feature values to ranks). This approach further improves the performance of the all-language model adaptation, even on average performing better than the native models. In more detail, the model outperforms a random classifier in more than 65% of low-resource settings, and even outperforms native models in 47% of these cases, while being on par with it in 21%. Notably, models trained solely on English fail to generalize well to low-resource languages, performing worse than random in over 70% of cases—highlighting the distinct editorial behaviors across linguistic communities.
Finally, since the size of the dataset increases with the combination of topics, we train the classifier with the maximum tree depth of 3 instead of 1, but do not find a substantial increase in performance.
Figure 6. Combined Model performances for different training datasets and strategies. Top panel: % of models that perform better, same as, or worse than a random classifier; Center: same, but compared to the native model; Bottom: violin plots of the corresponding distributions of the F1 macro performances. : ¡short description¿
4.4. In-Depth Examination of Source Usage
The above quantitative evaluation of the proposed model abstracts the rich insight that its output provides. When examining the domains that attain the highest reliability score by the classifier, we find interesting signals of the editorial treatments of different sources, in each topical sphere (here, we focus on English data). For instance, among the domains predicted (incorrectly) as reliable in the Climate Change domain is YouTube (which appeared in 644 articles in that dataset) with a positive (reliable) class probability of $P_{r}=0.64$, as well as Internet Movie Database (IMDB) (appearing in 56 articles, $P_{r}=0.56$). This points to the importance of popular or mainstream culture in this topic. When considering the History dataset, we find the classifier incorrectly labels as reliable domains pointing to sources that the perennial sources label as having partisan bias, such as the Jewish Virtual Library ($P_{r}=0.81$), Sixth Tone ($P_{r}=0.74$), and Daily Sabah ($P_{r}=0.69$).
On the other hand, one can use the output of this classifier to find sources that different communities of editors consider inappropriate for their topical domain. For instance, in the Biology dataset, Science-Based Medicine, which is considered generally reliable and has “a credible editorial board” according to the perennial source list, is treated by the editors more similarly to unreliable domains ($P_{r}=0.20$). Similarly, the Huffington Post, rated as “fairly reliable for factual reporting on non-political topics”, scores low in the COVID dataset ($P_{r}=0.07$), possibly due to the politically charged and serious nature of the rhetoric around the pandemic. As editorial policies of sources, as well as perceptions and behaviors of the Wikipedia editors, change, the metrics developed in this study may provide a guide for the improvement and updating of the perennial sources and potential adjustment of the labels in specific domains of expertise.
4.5. Alternative Indicators of Source Reliability
One limiting factor of our study was the availability of labels for the quality of sources. We consider an alternative: Media Bias / Fact Check (MBFC), a third-party website that provides ratings of “the bias, factual accuracy, and credibility of media sources” [^8] https://mediabiasfactcheck.com/about/, which has been widely used in misinformation research and by other indexes (Barret, 2021; Chowdhury et al., 2020; Ye and Skiena, 2019). We use this credibility metric and consider very high and high as the positive class, and low and very low as negative. With the same features and model as above and LOO cross-validation, we gauge its performance on MBFC. Although the class-average F1 macro is 0.61, the classifier performs especially poorly on the negative class, with a precision of 0.32 and recall of 0.55, suggesting that the features do not capture this metric. Indeed, there may be several reasons why such resources are not suitable for the analysis of Wikipedia references. First, there is little overlap between the sources noted in MBFC and in the English Wikipedia perennial sources list (out of 1156 perennial sources, 272 are in MBFC). Second, the MBFC credibility metric is weakly aligned with the perennial source reliability one. Third, MBFC itself is listed as generally unreliable by the English Wikipedia perennial sources list.
In conclusion, MBFC and similar resources are valuable for assessing source reliability in general research contexts. However, the above observations suggest that they are less suitable for evaluating the reliability of sources in the specific context of Wikipedia articles, especially if used as ground truth for supervised classification methods like those presented in this work.
5. Discussion
In this study, we have proposed and assessed editing behavior-based features for modeling the reliability of web domains used as references on Wikipedia. We show that these models can be a part of a language-agnostic toolkit applicable both to high- and low-resource languages. As such, this work extends previous efforts to use the edit history to measure the controversiality of Wikipedia articles (Borra et al., 2015). Editing behavior is not the only source of useful information for language-agnostic modeling. Links between articles (already used for topical clustering of Wikipedia articles (Johnson et al., 2021)) may provide semantically meaningful context for references, whereas users’ reading sessions (used for entity linking (Gerlach et al., 2021)) may give controversiality signals from the perspective of the audience. Future work could also take a page from the social media literature and build editor-resource networks (Shu et al., 2019) to capture structural peculiarities of topics, editors, and resources associated with low-quality content. Additionally, automated feature discovery methods, such as self-supervised or representation learning on structured edit data, could help uncover latent behavioral patterns not easily captured by manually designed features. However, this would come at the cost of losing interpretability, a key value of our current approach.
Apart from finding that the proposed features capture editing behavior within a dataset spanning one topic and language, we also observe that they retain useful information during the model adaptation across languages. Although English has been the lingua franca in NLP, having the best-developed tools and resources, we show that adaptation from the English-language dataset to other languages is largely inferior to using a model that combines information from many languages. This improvement is especially noticeable in performance on low-resource languages, where native models often perform no better than random baselines, due to insufficient revision and citation activity for effective learning. However, when training on data pooled from all languages and applying quantile normalization to account for resource differences, the model significantly outperforms random baselines and matches or outperforms native models in most low-resource cases. This strategy highlights a practical and scalable pathway to empower under-resourced Wikipedia communities, even in the absence of extensive local data or tooling.
These findings are aligned with limitations identified by Das et al. (2024), who suggest that, although language-agnostic features are used, limiting data selection to one language may limit the breadth of the captured behaviors and their applicability to new settings. As multilingual embedding models have gained popularity (Zhang et al., 2023a), it remains to be seen to what extent the latest developments in NLP will benefit the low-resource languages (Zhang et al., 2023b), and what the impact will be on Wikipedia specifically (Johnson et al., 2024). In this context, the need to monitor possible changes in the editorial standards and citation practices across languages becomes especially important. Community efforts could prioritize on developing robust cross-lingual transfer strategies, rather than relying solely on local data that may be too sparse to train reliable predictors
As ground truth, we have used the list of perennial sources (Wikipedia contributors, 2024) from the English Wikipedia, which is part of Wikipedia’s efforts to provide clear editorial instructions to its community. We compare this resource with another popular source of reliability measure, the MBFC (Zandt, 2024), and find significant differences between them, pointing to the necessity of Wikipedia-specific models like ours. Previous work comparing the perennial sources lists in different languages showed disagreement on whether some domains are indeed reliable (Baigutanova et al., 2023b). Therefore, the system we propose could be used to expand lists of perennial sources and lists derived from them [^9] https://science.feedback.org/consensus-credibility-scores-comprehensive-dataset-web-domains-credibility/, or even to assess cultural biases in Wikipedia sources (Zheng et al., 2023; Yang and Colavizza, 2024).
Limitations
The present study has multiple limitations. First, using language-agnostic approaches may be seen as a limitation, as it does not take advantage of the predictive power of language-dependent features. However, despite the significant advances with multilingual NLP resources such as mBERT (Devlin et al., 2019), the motivation for our model was precisely to be ready for use in the 300+ language editions of Wikipedia, especially those that are less privileged and under-resourced.
Second, while the dataset contains thousands of articles and cumulatively millions of revisions, it spans only a fraction of Wikipedia. Despite choosing topical foci such that they have different reference quality and time scales, they may not capture the entire diversity of subject matter of this massive collaborative effort. For instance, differentiating topics by whether the topic is currently controversial may reveal shifts in editorial handling of sources. Additionally, future versions of our modeling approach could incorporate newly generated data by implementing re-training strategies using updated activity data, including changes in the classification of web domain reliability scores from the perennial source list.
Third, our approach may inherently favor generalist sources over topic-specific ones. Because several behavioral features (e.g., the number of articles or revisions in which a source domain appears) capture overall prevalence within Wikipedia, broadly used domains tend to accumulate stronger permanence and visibility signals. Conversely, specialized domains, despite being authoritative within their fields, may appear in fewer contexts and thus receive lower reliability estimates. This limitation suggests a potential bias toward sources covering a wider topical range. Its presence could be investigated in future research and mitigated by normalizing the corresponding features by topical diversity.
Fourth, the selection of the English perennial sources list as the gold standard disadvantaged other languages. Unfortunately, only a few other languages have a similar resource, and they are significantly less extensive and conclusive (Baigutanova et al., 2023b). Given the global role played historically by the English edition of Wikipedia (Hale, 2014), our selection provided the best choice to reach the maximum potential coverage across other language editions. However, the adaptation of the domain list from English may not coincide with possibly different cultural or local interpretations of quality by the editors from other language communities. In such cases, it may be preferable to apply NLP transfer methods (Das et al., 2025) that leverage higher-resource language editions to support lower-resource ones, particularly when they are comparable in the domain of source reliability. Our study, which employs language-agnostic features, can serve as a foundation for identifying appropriate language pairs for transfer in future work.
Fifth, our model relies on a fixed ground truth, consistent with prior work (e.g., (Baigutanova et al., 2023a, b)). However, perennial source lists evolve over time as editorial consensus changes and new sources are evaluated. While we treat it as static for this study, future work should incorporate time-aware labels and explore retraining strategies to adapt to such changes and maintain alignment with current community standards.
Finally, we show that the method is sensitive to the amount of data available for training (and testing), possibly resulting in insignificant results in some low-resource language settings. Further research will be needed to extend the amount of available data in these contexts.
Ethical Considerations
This study, and online content moderation in general, have important ethical aspects that should be kept in mind by both the platforms and the misinformation researchers.
First, this study has a direct impact on the access equity to encyclopedic resources in low-resource languages. Currently, the lack of resources, such as the perennial source list for these versions of Wikipedia, limits the extent to which our model is able to reflect the standards of that linguistic community. The domain scores provided by the models adapted from other (higher resource) languages that we propose here could be a starting point for bootstrapping the editorial discussion of the quality of sources used in low-resource languages.
Second, the process of determining the quality of sources bears scrutiny. In the news domain, the quality of information is often determined by professional fact-checkers. However, on Wikipedia, the editors from the community define the status of the sources via a deliberation process, guided by the community standards. Any potential biases in the judgment of these editors or failings in the deliberative process would be reflected in the quality of these labels (for instance, according to Kharazian et al. (2024), some Wikipedia editions were dominated for a time by a “small group of administrators who introduced far-right bias and outright disinformation”). Generally, in 2014 Shaw and Hill (2014) concluded that the Wikipedia contributors self-organize into oligarchic structures, though in 2021 Rijshouwer et al. (2023) found “strong counter-tendencies” to editorial power concentration. Our proposed method may reveal community behaviors that are not necessarily explicitly stated, but which shape the platform’s use of outside sources, potentially aiding the community to self-monitor the use of sources.
We believe that the code and resources we make available to the community [^10] https://github.com/JacopoDignazi/Wiscom (compiled in accordance with FAIR (Findable, Accessible, Interoperable, and Reusable principles (Wilkinson et al., 2016)) will be useful to assist future research efforts and to further evaluate and extend the features and modeling we proposed.
Acknowledgements.
This work is partially supported by a grant from the Credibility Coalition, a project of Meedan and Hacks/Hackers. Authors also acknowledge support from the Lagrange Project of the Institute for Scientific Interchange Foundation (ISI Foundation), funded by Fondazione Cassa di Risparmio di Torino (Fondazione CRT). Earlier versions of this paper benefited from valuable feedback from Diego Sáez Trumper, Francisco Navas, Aitolkyn Baigutanova, and attendees of Wiki Workshop 2024.
Appendix A Appendix A: Feature list
We provide here a more detailed list of the 52 features used by our models.
A.1. Popularity features
  1. (1)
    $N_{\mathrm{articles}}$: Number of articles a domain has appeared in (sensitive to the size of the dataset).
  2. (2)
    $\overline{N_{articles}}$: Number of articles a domain has appeared in, normalized by the number of articles in the dataset.
  3. (3)
    $CurrN_{articles}$: Number of articles a domain is used in at collection time.
  4. (4)
    $\overline{CurrN_{articles}}$: Number of articles a domain is used in at collection time, normalized by the total number of articles in the dataset.
A.2. Permanence features
  1. (5)
    $\Sigma Perm_{d}$: Permanence, how long a domain has been used in an article (not necessarily consecutively), summed across all articles and measured in days.
  2. (6)
    $\Sigma Perm_{r}$: Permanence, how long a domain has been used in an article (not necessarily consecutively), summed across all articles and measured by his number of revisions.
  3. (7)
    $\Sigma CurrPerm_{d}$: How long a domain has been used in an article (not necessarily consecutively), summed across all articles and measured in days, but considering only the articles where the domain is currently used.
  4. (8)
    $\Sigma CurrPerm_{r}$: How long a domain has been used in an article (not necessarily consecutively), summed across all articles and measured in number of revisions, but considering only the articles where the domain is currently used.
  5. (9)
    $\overline{\Sigma Perm_{d}}$: Sum of all permanences for a domain in all articles, normalized by the sum of the ages of all articles, measured by the number of days.
  6. (10)
    $\overline{\Sigma Perm_{r}}$: Sum of all permanences for a domain in all articles, normalized by the sum of the ages of all articles, measured by the number of revisions.
  7. (11)
    $\overline{\Sigma CurrPerm_{d}}$: Sum of all permanences for a domain in articles where the domain was used at collection time, normalized by the sum of the ages of all articles, measured by the number of days.
  8. (12)
    $\overline{\Sigma CurrPerm_{r}}$: Sum of all permanences for a domain in articles where the domain was used at collection time, normalized by the sum of the ages of all articles, measured by the number of revisions.
  9. (13)
    $\langle Perm_{d}\rangle$: The average permanence of a domain (over all articles where it was used), measured by the number of days.
  10. (14)
    $\langle Perm_{r}\rangle$: The average permanence of a domain (over all articles where it was used), measured by the number of revisions.
  11. (15)
    $\langle SelfPerm_{d}\rangle$, : Self-permanence: permanence divided by the number of days since the first time the domains were added to the article, then averaged article-wise.
  12. (16)
    $\langle SelfPerm_{r}\rangle$: Self-permanence: permanence divided by the number of revisions since the first time the domains were added to the article, then averaged article-wise.
  13. (17)
    $\Sigma age_{d}$: The sum of ages of the domain over all the articles in a dataset, where the age of a domain in an article is the number of days since a URL from that domain has been first added to the article.
  14. (18)
    $\Sigma age_{r}$: The sum of ages of the domain over all the articles in a dataset, where the age of a domain in an article is the number of revisions since a URL from that domain has been first added to the article.
  15. (19)
    $\langle age_{d}\rangle$: The average age of a domain over all the articles it appears at least once, measured by the number of days.
  16. (20)
    $\langle age_{r}\rangle$: The average age of a domain over all the articles it appears at least once, measured by the number of revisions.
A.3. User-based features
  1. (21)
    $U_{add}$ Number of unique users that have added a domain.
  2. (22)
    $U_{start}$ Number of unique users that have added a domain for the first time on a page.
  3. (23)
    $U_{rem}$ Number of unique users that have removed a domain.
  4. (24)
    $U_{end}$ Number of unique users that have removed a domain for the last time on a page.
  5. (25)
    $R_{add}$ Number of unique registered users that have added a domain.
  6. (26)
    $R_{start}$ Number of unique registered users that have added a domain for the first time on a page.
  7. (27)
    $R_{rem}$ Number of unique registered users that have removed a domain.
  8. (28)
    $R_{end}$ Number of unique registered users that have removed a domain for the last time on a page.
  9. (29)
    $\overline{U_{add}}$ Number of unique users that have added a domain, divided by the total number of unique users in the dataset.
  10. (30)
    $\overline{U_{start}}$ Number of unique users that have added a domain for the first time on a page, divided by the total number of unique users in the dataset.
  11. (31)
    $\overline{U_{rem}}$: Number of unique users that have removed a domain, divided by the total number of unique users in the dataset.
  12. (32)
    $\overline{U_{end}}$: Number of unique users that have removed a domain for the last time on a page, divided by the total number of unique users in the dataset.
  13. (33)
    $\overline{R_{add}}$: Number of unique registered users that have added a domain, divided by the total number of unique users in the dataset.
  14. (34)
    $\overline{R_{start}}$: Number of unique registered users that have added a domain for the first time on a page, divided by the total number of unique users in the dataset.
  15. (35)
    $\overline{R_{rem}}$: Number of unique registered users that have removed a domain, divided by the total number of unique users in the dataset.
  16. (36)
    $\overline{R_{end}}$: Number of unique registered users that have removed a domain for the last time on a page, divided by the total number of unique users in the dataset.
  17. (37)
    $\langle{U_{add}}\rangle$: Average per page number of unique users that have added a domain.
  18. (38)
    $\langle{U_{start}}\rangle$: Average per page number of unique users that have added a domain for the first time on a page.
  19. (39)
    $\langle{U_{rem}}\rangle$: Average per page number of unique users that have removed a domain.
  20. (40)
    $\langle{U_{end}}\rangle$: Average per page number of unique users that have removed a domain for the last time on a page.
  21. (41)
    $\langle{R_{add}}\rangle$: Average per page number of unique registered users that have added a domain.
  22. (42)
    $\langle{R_{start}}\rangle$: Average per page number of unique registered users that have added a domain for the first time on a page.
  23. (43)
    $\langle{R_{rem}}\rangle$: Average per page number of unique registered users that have removed a domain
  24. (44)
    $\langle{R_{end}}\rangle$: Average per page number of unique registered users that have removed a domain for the last time on a page.
  25. (45)
    $Ratio(R_{add}/U_{add})$: The proportion of registered vs. all users that ever added a domain.
  26. (46)
    $Ratio(R_{start}/U_{start})$: The proportion of registered vs. all users that ever added a domain for the first time on a page.
  27. (47)
    $Ratio(R_{rem}/U_{rem})$: The proportion of registered vs. all users that ever removed a domain.
  28. (48)
    $Ratio(R_{end}/U_{end})$: The proportion of registered vs. all users that ever removed a domain for the last time on a page.
  29. (49)
    $Proba(R_{add})$: Probability that, when a domain is added, this is done by the registered users.
  30. (50)
    $Proba(R_{start})$: Probability that, when a domain is added for the first time on a page, this is done by the registered users.
  31. (51)
    $Proba(R_{rem})$: Probability that, when a domain is removed, this is done by the registered users.
  32. (52)
    $Proba(R_{end})$: Probability that, when a domain is removed for the last time on a page, this is done by the registered users.
Appendix B Appendix B: Medians of resource-wise statistics
Table A1. Statistics of the topic-language datasets used in the study grouped by high, mid, and low resource languages. All numbers are medians (apart from the number of languages). Perennial Domains (P. Domains) refers to the number of domains for the perennial source list in English found in the different datasets. The corresponding mean values can be found in Table 1.
Topic Langs Articles Revs Revs/Ar URLs URLs/Ar Domains P. Domains
High Climate change 7 1162 251085 62 34352 18 11992 191
COVID-19 7 595 136470 49 29826 21 6031 185
Biology 7 8228 388729 17 32315 3 8377 127
History 7 4839 417275 40 30957 5 9324 146
Media 7 3542 398138 46 43849 8 12441 240
Mid Climate change 36 325 23000 27 4399 9 1991 85
COVID-19 37 182 10026 19 4891 9 1326 99
Biology 37 2053 57234 11 5005 2 1643 59
History 37 1687 70988 26 5500 3 1901 62
Media 37 799 30098 20 5347 4 1991 123
Low Climate change 56 100 2403 12 633 6 373 25
COVID-19 75 24 518 9 466 9 214 31
Biology 41 440 $49$ 7 541 3 301 14
History 55 421 8020 13 449 3 256 16
Media 70 73 1063 6 263 3 148 22
Appendix C APPENDIX C: Language-specific dataset sizes
Table A2. Number of articles per language editions in the different datasets.
Language Climate change COVID-19 Biology History Media
de 479 274 463 549 472
en 975 856 931 1092 1460
es 339 258 397 461 509
fr 412 262 388 546 476
ja 217 145 230 246 382
pt 223 125 248 258 296
ru 257 162 317 435 377
ar 195 148 196 258 189
az 62 43 88 126 73
bg 109 44 141 175 116
bn 82 51 92 95 80
ca 176 94 217 242 148
cs 168 79 188 203 143
da 141 43 143 161 119
el 104 55 112 211 95
et 101 58 105 122 73
eu 104 36 94 144 68
fa 156 89 171 229 168
fi 184 68 184 177 178
he 184 146 230 301 225
hi 98 42 78 76 58
hr 99 33 104 148 83
hu 151 61 205 227 186
hy 56 39 70 112 83
id 136 83 129 154 122
it 259 161 289 426 427
ko 139 106 153 204 204
ms 73 63 82 91 67
nb 174 72 170 183 145
nl 238 114 270 295 233
pl 185 89 233 279 236
ro 106 45 122 171 109
simple 136 53 140 137 118
sk 105 26 118 137 91
sl 106 39 99 156 65
sr 115 59 129 230 126
sv 187 79 204 221 201
th 107 61 123 138 125
tr 157 89 160 247 182
uk 162 212 173 194 150
uz 36 39 33 84 42
vi 137 217 145 161 145
zh 205 246 212 250 326
$k$ 0 17 55 49 34
af 72 28 71 102 46
als 33 0 0 58 0
as 21 13 23 18 13
ast 84 20 88 130 84
ba 28 6 0 36 14
bcl 21 9 16 27 12
be 56 27 61 90 48
be-x-old 45 14 0 77 42
bs 74 21 71 118 52
ckb 31 16 0 44 38
cv 25 0 0 0 25
cy 63 23 556 161 55
dag 29 0 0 0 0
eo 104 26 121 137 79
ga 43 24 44 76 47
gl 108 34 127 152 81
gu 29 6 0 0 9
ha 32 23 0 20 8
ia 31 11 0 0 0
ig 17 0 7 10 9
jv 46 15 63 74 32
ka 66 24 76 113 77
kn 51 28 40 33 35
ku 30 13 0 58 31
la 69 19 89 130 54
lb 33 0 0 0 53
lt 102 26 120 129 79
lv 80 27 95 93 63
mk 72 18 73 124 59
ml 70 39 83 81 49
mn 43 0 42 0 18
mni 13 0 0 0 0
mnw 4 5 0 0 0
mr 63 20 0 0 33
my 37 18 0 29 16
ne 35 16 42 0 23
$n$ 87 14 83 101 59
oc 55 0 0 98 38
or 16 18 0 0 14
pa 34 24 29 36 25
pnb 28 9 38 38 16
ps 21 9 0 23 0
rm 6 0 0 0 0
rw 14 0 19 0 0
sh 67 20 62 100 52
si 51 16 32 24 16
sq 56 36 50 104 60
su 30 19 44 0 13
sw 44 14 47 74 31
ta 90 33 88 68 57
te 49 33 61 40 47
tg 30 9 0 0 23
tl 47 40 65 76 41
ur 57 37 58 99 42
war 34 0 0 48 17
zh-yue 58 54 62 59 49
am 0 6 0 0 0
ary 0 4 0 0 0
arz 0 23 32 79 38
awa 0 2 0 0 0
bjn 0 11 0 0 0
bo 0 4 0 0 0
br 0 13 0 90 41
bug 0 2 0 0 0
bxr 0 7 0 0 0
ceb 0 7 0 0 0
crh 0 3 0 0 0
dty 0 2 0 0 0
gn 0 6 0 0 0
gpe 0 1 0 0 0
ht 0 26 0 0 0
hyw 0 3 0 33 18
km 0 8 0 9 0
mg 0 7 0 0 0
min 0 17 0 0 0
pap 0 12 0 0 0
sco 0 13 31 31 24
shn 0 5 0 0 0
so 0 9 0 0 0
tk 0 6 0 0 0
ts 0 5 0 0 0
vo 0 3 0 0 12
wuu 0 15 0 0 0
xmf 0 10 0 0 17
yo 0 24 0 0 0
zh-min-nan 0 14 0 39 25
is 0 0 75 91 50
ky 0 0 21 20 12
pam 0 0 27 0 0
$t$ 0 0 26 57 14
azb 0 0 0 32 12
fo 0 0 0 26 16
fy 0 0 0 81 32
io 0 0 0 79 0
tum 0 0 0 2 0
an 0 0 0 0 35
frp 0 0 0 0 6
glk 0 0 0 0 3
lo 0 0 0 0 5
mai 0 0 0 0 8
nds 0 0 0 0 28
qu 0 0 0 0 15
sat 0 0 0 0 12
vec 0 0 0 0 19
Table A3. Number of revisions per language editions in the different datasets.
Language Climate change COVID-19 Biology History Media
de 344149 195448 388729 549521 398138
en 1516501 1036268 2392272 2213972 3850137
es 251085 142608 412920 417275 447902
fr 261336 136470 401974 570207 405992
ja 59235 37293 97684 117733 222585
pt 82137 32740 178778 153532 154532
ru 90492 56512 230846 348758 257507
ar 64147 52221 123245 164152 73088
az 5396 3275 37635 34400 6289
bg 14688 4606 50747 77809 24812
bn 9138 6567 12759 20227 7907
ca 43861 13878 219014 168055 45265
cs 37535 15008 70480 94004 36957
da 21730 4267 33759 54452 21325
el 15873 7175 16066 93809 12233
et 12000 6702 21939 28705 6906
eu 12238 2086 82623 45030 6865
fa 34394 14882 65428 89813 59550
fi 41846 10081 76296 70988 58117
he 45760 36839 84560 153974 74220
hi 10574 3708 7683 12406 4471
hr 11084 2842 21316 40023 9546
hu 27013 14401 93431 100103 50952
hy 5149 4177 10735 29745 12491
id 23373 14547 111305 55091 32762
it 102487 46926 204053 375945 347881
ko 23966 28724 57234 84491 88325
ms 6906 7698 11679 20204 7208
nb 36680 10026 81352 78627 40317
nl 78355 25913 303604 171942 93987
pl 45284 22540 160392 173581 116880
ro 14046 7347 25254 60565 19300
simple 30367 8317 31239 33583 23031
sk 11945 2913 18615 40322 9525
sl 12320 2265 12780 39772 5602
sr 14728 6756 46083 102224 20365
sv 49949 11079 264477 114709 71292
th 12084 11217 22152 36162 30098
tr 34878 19488 61404 121987 56558
uk 33040 55024 105302 102155 45483
uz 2067 3116 2961 17800 2682
vi 22628 66920 407504 45861 30883
zh 53587 131339 131500 122649 182721
$k$ 0 517 6564 5187 1523
af 6384 2827 12553 20980 3448
als 1320 0 0 4863 0
as 814 466 1025 590 270
ast 7812 997 40421 36452 8628
ba 1316 50 0 2004 449
bcl 1036 182 439 1082 158
be 3892 1473 10090 23364 2899
be-x-old 2385 518 0 12592 1722
bs 5029 1324 7127 30720 2982
ckb 1397 564 0 3609 2026
cv 860 0 0 0 746
cy 4720 1280 521785 50767 9683
dag 790 0 0 0 0
eo 11720 1920 29807 42486 7768
ga 2570 1830 4293 8020 2404
gl 13814 2608 32768 44564 11127
gu 1191 311 0 0 157
ha 1676 638 0 628 131
ia 1437 265 0 0 0
ig 884 0 116 284 166
jv 2325 492 $49$ 12716 1271
ka 5100 1312 11997 33511 8342
kn 2643 1366 2085 1482 1593
ku 1659 656 0 4968 1020
la 5306 749 11783 40788 3739
lb 1925 0 0 0 2999
lt 11336 $16$ 32199 34848 7025
lv 7180 4234 13517 12402 5800
mk 6096 811 8133 35526 4200
ml 6461 2373 14983 7912 3545
mn 2421 0 2011 0 474
mni 385 0 0 0 0
mnw 160 217 0 0 0
mr 4801 1140 0 0 1494
my 1469 1358 0 1317 430
ne 1383 487 3090 0 684
$n$ 8312 1801 10620 20234 4729
oc 4226 0 0 19691 1618
or 313 916 0 0 286
pa 1609 825 1498 2143 832
pnb 1076 100 4884 4351 290
ps 693 144 0 850 0
rm 148 0 0 0 0
rw 390 0 400 0 0
sh 4436 619 13554 28448 5352
si 2728 421 1249 869 469
sq 3919 2758 3169 24390 4022
su 1328 523 1885 0 389
sw 3057 416 2273 13290 1027
ta 10484 2321 12039 6974 4530
te 3599 3142 4649 2462 3206
tg $14$ 331 0 0 637
tl 3262 2685 6835 11473 2215
ur 4132 2055 3941 18574 2528
war 1436 0 0 7500 419
zh-yue 3844 6256 6081 7762 3438
am 0 115 0 0 0
ary 0 539 0 0 0
arz 0 1221 12611 18006 10561
awa 0 48 0 0 0
bjn 0 153 0 0 0
bo 0 57 0 0 0
br 0 530 0 22230 2022
bug 0 7 0 0 0
bxr 0 238 0 0 0
ceb 0 125 0 0 0
crh 0 50 0 0 0
dty 0 16 0 0 0
gn 0 89 0 0 0
gpe 0 31 0 0 0
ht 0 1304 0 0 0
hyw 0 118 0 1401 401
km 0 464 0 134 0
mg 0 138 0 0 0
min 0 518 0 0 0
pap 0 387 0 0 0
sco 0 223 1636 1609 689
shn 0 143 0 0 0
so 0 164 0 0 0
tk 0 95 0 0 0
ts 0 98 0 0 0
vo 0 33 0 0 412
wuu 0 341 0 0 0
xmf 0 370 0 0 342
yo 0 1353 0 0 0
zh-min-nan 0 393 0 3607 889
is 0 0 7778 12301 2541
ky 0 0 1317 705 297
pam 0 0 798 0 0
$t$ 0 0 2568 11215 459
azb 0 0 0 3732 1496
fo 0 0 0 1252 307
fy 0 0 0 12889 1099
io 0 0 0 15974 0
tum 0 0 0 16 0
an 0 0 0 0 1325
frp 0 0 0 0 36
glk 0 0 0 0 61
lo 0 0 0 0 47
mai 0 0 0 0 104
nds 0 0 0 0 801
qu 0 0 0 0 435
sat 0 0 0 0 133
vec 0 0 0 0 569